OpenStudy (anonymous):

Solve the system, using substitution. One point for correct steps, one point for correct ordered pair, (a,b). a + b = 4 and 2a + 3b = 11

OpenStudy (anonymous):

b = 4 - a So then 2a + 3(4-a) = 11 2a + 12 - 3a = 11 -a + 12 = 11 -a = -1 a = 1 Solve for b the same way. Hope this helps.

OpenStudy (anonymous):

thank you!:)